Complete the square x^2+8x-5=15
Answer by jim_thompson5910(35256)

x^2+8x-5=15

x^2+8x-5+5=15+5

x^2 + 8x = 20

x^2 + 8x + 16 = 20 + 16 ... take half of 8, then square it, add it to both sides

x^2 + 8x + 16 = 36


I'll let you finish up